# Enhancing Fairness in Unsupervised Graph Anomaly Detection through Disentanglement

### v1

@@ 3.4. Training and Inference
The overall training process of DEFEND is presented in Algorithm 1. In the disentangled representation learning phase, given an attributed graph $\mathcal{G}$, we aim to train a variational graph autoencoder, in which the disentangled encoder $f_{e}$ can separate nonsensitive representations while maintaining information for reconstructing $\mathcal{G}$ by the decoder $f_{a}$ and $f_{x}$. We encourage the disentanglement with a learnable adversary $g_{\omega}$, which is optimized with $\mathcal{L}_{adv}$. The optimization of $f_{e}$, $f_{a}$ and $f_{x}$ and the optimization of $g_{\omega}$ are conducted adversarially. For anomaly detection, we freeze the encoder $f_{e}$ to obtain sensitive-irrelevant node representations and reconstruct attributes solely with the decoder $f_{\phi}$. We utilize reconstruction error as the anomaly score and constrain the correlation between reconstruction error and predicted sensitive attributes.
We analyze the computational complexity of DEFEND as follows. In the disentangled representation learning phase, the encoder $f_{e}$ processes with a complexity of $O(E\times(d+m))$, where $E$ denotes the number of edges and $d+m$ denotes the attribute dimensionality. The decoders $f_{a}$ and $f_{x}$ exhibit a complexity of approximately $O(N^{2})$ due to adjacency matrix reconstruction. The adversary operates at $O(D_{0})$ with $D_{0}$ neurons in the MLP. In the anomaly detection phase, the complexity for the frozen encoder remains $O(E\times(d+m))$, whereas the MLP-based decoder $f_{\phi}$ exhibits a complexity of $O(D_{1})$ with $D_{1}$ neurons in the MLP. Given that $D_{0}\leq d$ and $D_{1}\leq d$ in most cases, the overall computational complexity is estimated as $O(N^{2}+E\times(d+m))$. Notably, in practical applications, the quadratic complexity associated with the matrix reconstruction is effectively reduced using sparse matrix techniques, which are well-suited for the sparse graphs commonly encountered in real-world scenarios.
@@
6:     Calculate $\mathcal{L}_{total}$ according to Eq. (missing) 15;
@@
8:     Calculate $\mathcal{L}_{adv}$ according to Eq. (missing) 16;
@@
11:
12:  // Anomaly detection phase;
13:  Freeze $f_{e}$ and initialize $f_{\phi}$;
14:  for $t=1,2,\dots,T$ do
15:     $\bar{\mathbf{Z}}_{x},\mathbf{Z}_{s}\leftarrow f_{e}(\mathbf{X},\mathbf{A})$;
16:     $\tilde{\mathbf{X}}\leftarrow f_{\phi}(\bar{\mathbf{Z}}_{x},\tilde{\mathbf{Z}}_{s})$;
17:     Calculate anomaly score $\mathbf{o}$ according to Eq. (missing) 17;
18:     Calculate $\mathcal{L}_{corr}$ according to Eq. (missing) 18;
19:     Calculate $\mathcal{L}_{ad}$ according to Eq. (missing) 19;
20:     Update $f_{\phi}$ by gradient descent using $\mathcal{L}_{ad}$;
21:  end for
22:  return  $\mathbf{o}$;
3.5. Discussion
We analyze the generalization of DEFEND on various types of sensitive attributes. (1) Multiple Binary Sensitive Attributes. Let $\mathbf{s}=\left[{s^{0}},{s^{1}},\ldots,{s^{m}}\right]$ denotes the sensitive attributes for each node, where $s^{j}\in\{0,1\}(0\leq j<m)$. The disentanglement between sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}\in\mathbb{R}^{N\times m}$ requires that $\mathbf{Z}_{x}$ is independent of each sensitive attribute $\mathbf{Z}_{s}^{(j)}$. Thus, $\mathbf{Z}_{x}$ and $\mathbf{Z}_{s}$ is disentangled if the aggregate posterior distribution can be factorized as $p(\mathbf{Z}_{x},\mathbf{Z}_{s})=p(\mathbf{Z}_{x})\prod_{j}p(\mathbf{Z}_{s}^{(j)})$. (2) Single Categorical/Continuous Sensitive Attributes. With single binary sensitive attributes, the conditional distribution $p(\mathbf{S}|\mathbf{Z}_{s})$ is typically modeled using a Bernoulli distribution. However, with categorical attributes, this shifts to a Multinomial distribution, i.e., $p(\mathbf{S}|\mathbf{Z}_{s})=\text{Multinomial}(\mathbf{S}|\text{SoftMax}(\mathbf{Z}_{s})).$ Similarly, for continuous attributes, $p(\mathbf{S}|\mathbf{Z}_{s})$ can be modeled using a Gaussian distribution.
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets.
We employ two real-world datasets for fair graph anomaly detection in our experiments, including Reddit and Twitter (Neo et al., 2024). Details of these datasets are summarized in Table 1.
  • Reddit: The Reddit dataset contains 110 politically oriented subreddits, encompassing all historical postings within these forums. It also includes all historical posts from several active discussion participants. A relational graph was constructed by linking users who posted in the same subreddit within 24 hours.
  • Twitter: The Twitter dataset is conducted on 47,712 users with their historical posts, user profiles, and follower relationships. The user information, such as the organization status, was inferred using the M3 System from user profiles and tweets. The key account metrics and averaged post embeddings were combined to form node features.

### v2

@@ Appendix A Pseudo Code
The overall training process of DEFEND is presented in Algorithm 1. Firstly, given an attributed graph $\mathcal{G}$, we aim to train a variational graph autoencoder, in which the disentangled encoder $f_{e}$ can separate sensitive-irrelevant representations while maintaining information for reconstructing $\mathcal{G}$ by the decoder $f_{a}$ and $f_{x}$ (Line 4-6). We encourage the disentanglement with a learnable adversary $g_{\omega}$ optimized by $\mathcal{L}_{adv}$ (Line 8-9). The optimization of $f_{e}$, $f_{a}$ and $f_{x}$ and that of $g_{\omega}$ are conducted adversarially. Next, the frozen $f_{e}$ captures sensitive-irrelevant node representations (Line 14), and the decoder $f_{\phi}$ solely reconstructs attributes (Line 15). We utilize reconstruction error as the anomaly score and constrain the correlation between reconstruction error and predicted sensitive attributes (Lines 16-19). Finally, only $f_{\phi}$ will be updated by $\mathcal{L}_{ad}$ in the anomaly detection phase.
@@
6:     Calculate $\mathcal{L}_{total}$ according to Equation 15;
@@
8:     Calculate $\mathcal{L}_{adv}$ according to Equation 16;
@@
11:  // Anomaly detection phase;
12:  Freeze $f_{e}$ and initialize $f_{\phi}$;
13:  for $t=1,2,\dots,T$ do
14:     $\bar{\mathbf{Z}}_{x},\mathbf{Z}_{s}\leftarrow f_{e}(\mathbf{X},\mathbf{A})$;
15:     $\tilde{\mathbf{X}}\leftarrow f_{\phi}(\bar{\mathbf{Z}}_{x},\tilde{\mathbf{Z}}_{s})$;
16:     Calculate anomaly score $\mathbf{o}$ according to Equation 17;
17:     Calculate $\mathcal{L}_{ad}$ according to Equation 19;
18:     Update $f_{\phi}$ by gradient descent using $\mathcal{L}_{ad}$;
19:  end for
20:  return  $\mathbf{o}$;
Appendix B Complexity Analysis
We analyze the time and space computational complexity of DEFEND. Let $N$, $E$, $d$ and $h$ denote the number of nodes, edges, attribute dimensions and hidden dimensions, respectively.
In the disentangled representation learning phase, the encoder $f_{e}$ exhibits a complexity of $O(Ed)$, while the decoders $f_{a}$ and $f_{x}$ exhibit a complexity of $O(N^{2})$ for adjacency matrix reconstruction. The adversary $g_{\omega}$ operates at $O(Nh)$. For loss computations, the variational reconstruction loss $\mathcal{L}_{vae}$ exhibits $O(N^{2})$ complexity for adjacency matrix reconstruction and $O(Nd)$ for attribute reconstruction. The disentanglement loss $\mathcal{L}_{dis}$, predictive loss $\mathcal{L}_{pre}$, and adversarial loss $\mathcal{L}_{adv}$ each exhibit $O(Nh)$ complexity. In the anomaly detection phase, the frozen encoder $f_{e}$ maintains $O(Ed)$ complexity, while the MLP-based decoder $f_{\phi}$ exhibits $O(Nh)$ complexity. Both the correlation constraint loss $\mathcal{L}_{corr}$ and reconstruction loss $\mathcal{L}_{rec}^{X}$ exhibit $O(Nh)$ complexity. Thus, the total time complexity is $O(N^{2}+Ed)$. The space complexity of DEFEND is dominated by the storage of the dense adjacency matrix $O(N^{2})$, node feature matrix $O(Nd)$, and model parameters with intermediate results $O(Nh)$, resulting in a total space complexity of $O(N^{2}+Nd+Nh)$.
To enhance scalability, we implement a sparse version that computes reconstruction only for existing edges and approximates the global statistics using row/column means, instead of reconstructing the full adjacency matrix. We also replace dense matrix operations with sparse operations in loss computations. This optimization reduces the complexity of the decoder $f_{a}$ to $O(Eh)$ for structure reconstruction. Besides, the complexity of $\mathcal{L}_{vae}$ decreases to $O(E)$ for structure and maintains $O(Nd)$ for attribute reconstruction. Thus, the overall time complexity is reduced to $O(Ed+Nd)$. The space complexity decreases significantly to $O(E+Nd+Nh)$ by utilizing sparse matrix representations and eliminating dense adjacency matrices.
Appendix C Datasets